# Diffusion Models for Time Series Forecasting: A Survey

### v1

@@ Abstract
Diffusion models, initially developed for image synthesis, demonstrate remarkable generative capabilities. Recently, their application has expanded to time series forecasting (TSF), yielding promising results. In this survey, we firstly introduce the standard diffusion models and their prevalent variants, explaining their adaptation to TSF tasks. We then provide a comprehensive review of diffusion models for TSF, paying special attention to the sources of conditional information and the mechanisms for integrating this conditioning within the models. In analyzing existing approaches using diffusion models for TSF, we provide a systematic categorization and a comprehensive summary of them in this survey. Furthermore, we examine several foundational diffusion models applied to TSF, alongside commonly used datasets and evaluation metrics. Finally, we discuss current limitations in these approaches and potential future research directions. Overall, this survey details recent progress and future prospects for diffusion models in TSF, serving as a reference for researchers in the field. [^1] Resources are available at https://github.com/synlp/TSF-Diff-Review.
Keywords
@@ 1 Introduction
Time series forecasting (TSF) is a fundamental task that aims to predict future values of a sequence based on historical data. It plays a vital role in a wide range of real-world applications, such as energy demand forecasting [1, 2, 3], traffic flow prediction [4, 5, 6], and healthcare monitoring [7, 8, 9]. Time series data present diverse structures, ranging from univariate sequences that monitor a single variable over time, to multivariate sequences capturing interactions among multiple variables, as well as more complex forms such as spatio-temporal data from geo-distributed sensors and multimodal sequences that combine temporal signals with text, images, or metadata. TSF approaches undergo a significant evolution, progressing from conventional statistical models such as autoregressive integrated moving average model (ARIMA) [10, 11] and exponential smoothing [12, 13, 14] to deep learning approaches such as recurrent neural networks (RNNs) [15, 16, 17], convolutional neural networks (CNNs) [18, 19, 20], and Transformers [21, 22, 23], where the deep learning based models are widely used in nowadays research.
@@
Among different deep learning approaches, generative ones rise in recent years with various representative models, including variational autoencoders (VAEs) [24, 25, 26], generative adversarial networks (GANs) [27, 28, 29], and especially diffusion models [30, 31, 32, 33, 34] attract growing interest for their ability to capture uncertainty and produce high-quality forecasts. Particularly, diffusion models, as latent variable generators, learn complex data distributions through a two-step cycle of progressive noising followed by reverse denoising. This paradigm revolutionizes computer vision, enabling high-fidelity image synthesis, inpainting, and super-resolution without the adversarial instability of GANs. Motivated by such success, diffusive modeling migrates from vision to other scenarios, including TSF, and provides a generative framework for capturing complex temporal patterns.
@@
Diffusion models for TSF are commonly formulated as conditional generative models, where the objective is to generate future sequences conditioned on past observations and, optionally, additional contextual information [26, 35, 36, 37]. During training, Gaussian noise is gradually added to the ground-truth future sequence through a forward diffusion process. A neural network guided by the historical context is then trained to reverse this noising process. At inference time, the model generates forecasts by iteratively denoising the noise, effectively sampling from the learned conditional distribution. Compared to conventional models, diffusion-based approaches offer several key advantages: (1) they produce probabilistic forecasts, enabling uncertainty quantification by sampling multiple plausible futures; (2) they exhibit stable training behavior and avoid adversarial instability; and (3) they offer flexible generative capabilities, allowing seamless integration with various conditioning signals such as covariates, control variables, or multimodal inputs. Therefore, it is essential to provide a comprehensive review of the diffusion models for TSF, given that most existing literature reviews of the TSF approaches focus on other approaches [38, 39, 40].
@@
In this paper, we provide a comprehensive survey of diffusion models for TSF. We begin by reviewing the foundation diffusion models, including Denoising Diffusion Probabilistic Model (DDPM) [32], Denoising diffusion implicit model (DDIM) [41], and their extensions incorporating Classifier Guidance [42] and Classifier-Free Guidance [43] (Section 2). We then propose a systematic taxonomy to classify existing diffusion-based TSF approaches along two dimensions: conditioning source (e.g., historical time series and multimodal data) (Section 3), and condition integration process (e.g., feature-centric condition and diffusion-centric condition) (Section 4). For each category, we summarize its common traits and highlight representative approaches. Furthermore, we review widely used datasets (Section 5) and evaluation metrics (Section 6). Finally, we present an overall analysis of existing approaches, highlight their current limitations (Section 7), and propose potential directions for future research (Section 8). Through this survey, we aim to synthesize recent advances, reveal design patterns, and provide insights into the evolving landscape of diffusion-based TSF.
Figure 1: General approach of diffusion-based TSF.
@@ 2 Foundation Diffusion Models for TSF
Standard diffusion models define a generative process that progressively adds noise to data through a forward trajectory and reconstructs the original sample via a reverse denoising process. Conditional diffusion models [^2] Unless otherwise specified, the “diffusion model” (as well as its specific version, e.g., DDPM [32]) in the following text stand for the conditional diffusion model. extend this framework by introducing auxiliary information, such as labels or historical time series, to guide the generation process. A representative diffusion model is DDPM [32], which achieves high-quality generation but requires many iterative sampling steps, resulting in slow inference speed. DDIM [41] addresses this limitation by reformulating the reverse process as a deterministic trajectory, which accelerates sampling while maintaining compatibility with the original model. In TSF, the objective is to generate future sequences conditioned on past observations, making the task a natural fit for conditional diffusion models. The conditioning input typically consists of historical time series, and the model learns to reconstruct future values from noise under this context. Recent research [44, 45, 46] in diffusion-based TSF also introduces guidance techniques to control the influence of conditioning signals during generation, enabling more flexible and robust forecasting. The following sections provide a detailed roadmap of these concepts, including the mathematical formulation of DDPM and DDIM, the adaptation of conditional diffusion models for TSF, and the integration of guidance mechanisms for enhanced controllability.
@@ 2.1 Denoising Diffusion Probabilistic Model (DDPM)
Denoising Diffusion Probabilistic Models (DDPMs) [32] define a generative process by progressively adding and removing noise through a sequence of forward and reverse steps, conditioned on auxiliary information. Let $\mathbf{Y}_{0}$ denote the original data sample, and $\mathbf{c}$ the conditioning input. DDPM constructs a forward diffusion process that corrupts $\mathbf{Y}_{0}$ into a noise distribution and a reverse process that reconstructs $\mathbf{Y}_{0}$ from noise conditioned on $\mathbf{c}$. The following illustrates the details of the forward diffusion process and the revserse denoising process.
Forward Diffusion Process
The forward process corrupts the data $\mathbf{Y}_{0}$ using a predefined noise variance schedule $\{\beta_{t}\}_{t=1}^{T}$, where $\beta_{t}$ is the noise variance at diffusion step $t$, and $T$ is the total diffusion steps. The entire forward process forms a Markov chain that is fomulated by
@@
where $q(\cdot\mid\cdot)$ denotes the forward process distribution and $\mathbf{Y}_{t}$ is the latent variable at diffusion step $t$. In addition, $\mathbf{Y}_{1:T}$ is a noisy data sequence from diffusion step $1$ to $T$. Meanwhile, $q(\mathbf{Y}_{t}\mid\mathbf{Y}_{t-1})=\mathcal{N}(\cdot;\boldsymbol{\mu},\boldsymbol{\Sigma})$ is a multivariate Gaussian distribution with mean $\boldsymbol{\mu}=\sqrt{1-\beta_{t}}\,\mathbf{Y}_{t-1}$ and covariance $\boldsymbol{\Sigma}=\beta_{t}\mathbf{I}$ (where $\mathbf{I}$ denotes the identity matrix). By simplification, the latent variable $\mathbf{Y}_{t}$ are sampled from a distribution directly conditioned on $\mathbf{Y}_{0}$:
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. The distribution of the endpoint $\mathbf{Y}_{T}$ of the forward process obeys a Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ with mean $\mathbf{0}$ and covariance $\mathbf{I}$ (i.e., $\mathbf{Y}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$).
Reverse Denoising Process
Contrary to the forward process, the reverse process reconstructs $\mathbf{Y}_{t-1}$ from $\mathbf{Y}_{t}$ via a transition distribution $p_{\theta}(\cdot\mid\cdot)$ parametrized by $\theta$, which is formulated as
$p_{\theta}(\mathbf{Y}_{t-1}\mid\mathbf{Y}_{t},\mathbf{c})=\mathcal{N}(\mathbf{Y}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{Y}_{t},\mathbf{c},t),\Sigma_{\theta}(\mathbf{Y}_{t},\mathbf{c},t)),$ (3)
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the Gaussian noise added in the forward process, and $\boldsymbol{\epsilon}_{\theta}(\cdot)$ is the noise prediction network with conditional $\mathbf{c}$ as input. In addition, $\mathbb{E}[\cdot]$ denotes the expectation operation and $\left\|\cdot\right\|^{2}$ represents Euclidean norm function. During inference, the model iteratively applies the conditional reverse transitions starting from $\mathbf{Y}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ as
$\mathbf{Y}_{t-1}\sim p_{\theta}(\mathbf{Y}_{t-1}\mid\mathbf{Y}_{t},\mathbf{c}),\quad t=T,T-1,\dots,1.$ (5)
@@ 2.2 Denoising Diffusion Implicit Model (DDIM)
Compared to DDPMs, DDIMs eliminates stochasticity in the reverse process by removing the random noise term at each step, while preserving the marginal distributions of the original diffusion trajectory. The step-skip sampling in DDIMs eliminates the need to iterate over all $T$ time steps in the reverse denoising process. Instead, DDIM samples only on a subset of diffusion steps $t_{s_{1}},t_{s_{2}},\dots,t_{s_{S}}$, where $t_{s_{1}}=T$ and $t_{s_{S}}=0$. Specifically, for sampling step $k$ from $1$ to $S-1$, the DDIMs computes $\mathbf{Y}_{t_{s_{k-1}}}$ from $\mathbf{Y}_{t_{s_{k}}}$ using Eq. (6) and replaces $t$ and $t-1$ with $t_{s_{k}}$ and $t_{s_{k-1}}$, respectively.
Beyond DDIMs, several advanced sampling techniques are proposed to accelerate diffusion models. For instance, Analytic-DPM [47] leverages closed-form solutions to the reverse diffusion ordinary differential equation, enabling direct computation of sample trajectories without iterative simulation. DPM-Solver [48] applies high-order numerical solvers to the reverse process, replacing the standard Euler approach with more accurate and efficient integration schemes. These approaches further reduce inference time and improve sample quality by optimizing the reverse trajectory computation.
@@ 2.3 Applying Diffusion Models for TSF
A straightforward approach to applying diffusion models to TSF is to set the original data $\mathbf{Y}_{0}$ as the future time series $\mathbf{x}_{1:H}=(x_{1},x_{2},\ldots,x_{i},\ldots x_{H})$, where $H$ is the prediction horizon and $x_{i}$ denotes the time series value at timestep $i$. Meanwhile, the condition $\mathbf{c}$ in diffusion model equals to the historical time series $\mathbf{x}_{-L+1:0}=(x_{-L+1},x_{-L+2},\ldots,x_{0})$ with $L$ denoting the length of the history time series. In this formulation, the model learns the conditional distribution $q(\mathbf{x}_{1:H}\mid\mathbf{x}_{-L+1:0})$ of future values given historical time series $\mathbf{x}_{-L+1:0}$. The forward diffusion process corrupts the future time series $\mathbf{x}_{1:H}$ with Gaussian noise over $T$ steps, while the reverse process reconstructs $\mathbf{x}_{1:H}$ from noise $\mathbf{x}^{T}_{1:H}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, conditioned on the historical time series $\mathbf{x}_{-L+1:0}$. Specifically, the DDPM for TSF, as illustrated in Figure 1, defines the forward process as
@@
Meanwhile, the reverse process is parameterized as
$p_{\theta}(\mathbf{x}_{1:H}^{t-1}\mid\mathbf{x}_{1:H}^{t},\mathbf{x}_{-L+1:0})=\mathcal{N}(\mathbf{x}_{1:H}^{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{1:H}^{t},\mathbf{x}_{-L+1:0},t),\Sigma_{\theta}(\mathbf{x}_{1:H}^{t},\mathbf{x}_{-L+1:0},t))$ (10)
@@
The DDIM variant for TSF also modifies the reverse process to a deterministic trajectory which is formulated as
@@
In the field of conditional generation for vision, guidance techniques are proposed to flexibly control the strength of conditioning and improve sample quality. Recent studies in diffusion-based TSF also adapt guidance mechanisms to incorporate additional conditioning sources without modifying model architectures. Guidance mechanisms provide a way to steer the generation process toward desired properties or constraints. One of these techniques, named classifier guidance, modifies the reverse process by adding the gradient of a classifier [^3] Although TSF denotes this component as a predictor, we use the term “classifier” here to remain consistent with the original technique’s nomenclature. $p_{\text{clf}}(\mathbf{x}_{-L+1:0})$ trained to predict the historical time series $\mathbf{x}_{-L+1:0}$ from the noisy future sequence $\mathbf{x}_{1:H}^{t}$. According to Dhariwal and Nichol [42], the guidance is simplified to an update to the mean function:
@@
Different from the classifier guidance that requires additional training of a classifier, classifier-free guidance [43] trains the diffusion model to handle both conditional and unconditional generations by randomly dropping the conditioning input during training. At inference, the model combines the conditional and unconditional predictions to update the predicted noise:
@@
where $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{1:H}^{t},\varnothing,t)$ is the unconditional noise prediction and $\varnothing$ represents the historical time series is set to a null tensor. These mechanisms enable flexible integration of historical and auxiliary information in diffusion-based TSF models.

### v2

@@ Abstract
Diffusion models, initially developed for image synthesis, demonstrate remarkable generative capabilities. Recently, their application has expanded to time series forecasting (TSF), yielding promising results. Existing surveys on time series primarily focus on the application of diffusion models to time series tasks or merely provide model-by-model introductions of diffusion-based TSF models, without establishing a systematic taxonomy for existing diffusion-based TSF models. In this survey, we firstly introduce several standard diffusion models and their prevalent variants, explaining their adaptation to TSF tasks. Then, we provide a comprehensive review of diffusion models for TSF, paying special attention to the sources of conditional information and the mechanisms for integrating this conditioning within the models. In analyzing existing approaches using diffusion models for TSF, we provide a systematic categorization and a comprehensive summary of them in this survey. Furthermore, we examine several foundational diffusion models applied to TSF, alongside commonly used datasets and evaluation metrics. Finally, we discuss the progress and limitations of these approaches, as well as potential future research directions for diffusion-based TSF. Overall, this survey offers a comprehensive overview of recent progress and future prospects for diffusion models in TSF, serving as a valuable reference for researchers in the field.
Index Terms:
@@
22footnotetext: Work done at University of Science and Technology of China.
@@ 1 Introduction
Time series forecasting (TSF) is a fundamental task that aims to predict future values of a sequence based on historical data. It plays a vital role in a wide range of real-world applications, such as energy demand forecasting [1, 2, 3], traffic flow prediction [4, 5, 6], and healthcare monitoring [7, 8, 9]. Time series data present diverse structures, ranging from univariate sequences that monitor a single variable over time [10, 11, 12], to multivariate sequences capturing interactions among multiple variables [13, 14, 15], as well as more complex forms such as spatio-temporal data from geo-distributed sensors [16, 17, 18] and multimodal sequences that combine temporal signals with text, images, or metadata [19, 20, 21]. TSF approaches undergo significant evolution. Early approaches primarily rely on conventional statistical models, such as the autoregressive integrated moving average (ARIMA) model [22, 23] and exponential smoothing techniques [24, 25, 26]. Subsequently, the field shifts towards deep learning-based models. These include recurrent neural networks (RNNs) [27, 28, 29, 30], convolutional neural networks (CNNs) [31, 32, 33, 34], and Transformers [35, 36, 37, 38, 39]. More recently, large-scale Transformer-based large language models (LLMs) [40, 41, 42] and foundation models [43, 44, 45] demonstrate powerful zero-shot learning capabilities for TSF and achieve remarkable performance.
@@
Among different deep learning approaches, generative ones rise in recent years with various representative models, including variational autoencoders (VAEs) [46, 47, 48], generative adversarial networks (GANs) [49, 50, 51], and especially diffusion models [52, 53, 54, 55] that attract growing interest for their ability to capture uncertainty and produce high-quality predictions. Particularly, diffusion models, as latent variable generators, learn complex data distributions through a two-step cycle of progressive noising followed by reverse denoising. This paradigm advances computer vision, enabling high-fidelity image synthesis, inpainting, and super-resolution without the adversarial instability of GANs [56, 57]. Motivated by such success, diffusive modeling migrates from vision to a wide range of data modalities [58, 59, 60, 61], including the time series data [62, 63] and provides a robust generative framework for capturing complex temporal patterns.
@@
Diffusion models for TSF are commonly formulated as conditional generative models, where the objective is to generate future sequences conditioned on past observations and, optionally, additional contextual information [48, 62, 64]. During training, Gaussian noise is gradually added to the ground-truth future sequence through a forward diffusion process. A neural network guided by the historical context is then trained to reverse this noising process. At inference time, the model generates forecasts by iteratively denoising the noise, effectively sampling from the learned conditional distribution. Compared to conventional models, diffusion-based approaches offer several key advantages: (1) they produce probabilistic forecasts, enabling uncertainty quantification by sampling multiple plausible futures; (2) they exhibit stable training behavior and avoid adversarial instability; and (3) they offer flexible generative capabilities, allowing seamless integration with various conditioning signals such as covariates, control variables, or multimodal inputs [65, 66, 67]. Critically, diffusion models demonstrate strong empirical performance, achieving competitive or state-of-the-art performance on various TSF benchmark datasets [68, 67, 69]. Consequently, diffusion models are attracting increasing attention within the TSF research community. Therefore, it is essential to provide a comprehensive review of the diffusion models for TSF, given that most existing literature reviews of the TSF approaches focus on other approaches [70, 71, 72, 73]. Furthermore, existing surveys on diffusion models in time series applications [74, 75, 76] focus more on how diffusion models are applied to various time series tasks, or provide a model-by-model description of existing diffusion-based TSF studies. Consequently, a dedicated review is needed that offers a systematic and hierarchical taxonomy to organize and synthesize the current landscape of diffusion-based TSF research.
@@
In this survey, we provide a comprehensive survey of diffusion models for TSF. We begin by reviewing the foundation diffusion models, including Denoising Diffusion Probabilistic Model (DDPM) [54], Denoising diffusion implicit model (DDIM) [77], and their adaptation for TSF tasks incorporating Classifier Guidance [78] and Classifier-Free Guidance [79] (Section 2). We then propose a systematic taxonomy to classify existing diffusion-based TSF approaches along two dimensions: conditioning source (e.g., historical time series and multimodal data) (Section 3), and condition integration process (e.g., feature-centric condition and diffusion-centric condition) (Section 4). For each category, we summarize its common traits and highlight representative approaches. Furthermore, we review widely used datasets (Section 5) and evaluation metrics (Section 6). Finally, we present an overall analysis of existing approaches, highlight their current progress and limitations (Section 7), and propose potential directions for future research (Section 8). Through this survey, we aim to comprehensively synthesize recent advances, reveal design patterns, and provide deeper insights into the evolving landscape of diffusion-based TSF.
Figure 1: General framework of diffusion-based TSF. The left side illustrates the division of time series into a historical window (past time steps) and a future horizon (target sequence). The top-right presents the diffusion process: the forward diffusion process progressively adds Gaussian noise to the clean future time series, while the reverse denoising process reconstructs the future sequence. The bottom-right details the time series conditional diffusion model’s core components: Condition Source (Historical Time Series and Multi-modal Data) and Condition Integration (Feature-centric and Diffusion-centric). The black arrows denote the forward diffusion process, and the red arrows represent the model’s inference process. Historical time series are preprocessed or combined with multi-modal data to form the condition source, which is then integrated via feature-centric denoising networks or diffusion-centric specialized diffusion processes.
@@ 2 Foundation Diffusion Models for TSF
Standard diffusion models define a generative process that progressively adds noise to data through a forward trajectory and reconstructs the original sample via a reverse denoising process. Conditional diffusion models [^1] Unless otherwise specified, the “diffusion model” (as well as its specific version, e.g., DDPM [54]) in the following text stand for the conditional diffusion model. extend this framework by introducing auxiliary information, such as labels or historical time series, to guide the generation process. A representative diffusion model is DDPM [54], which achieves high-quality generation but requires many iterative sampling steps, resulting in slow inference speed. DDIM [77] addresses this limitation by reformulating the reverse process as a deterministic trajectory, which accelerates sampling while maintaining compatibility with the original model. In TSF, the objective is to generate future sequences conditioned on past observations, making the task a natural fit for conditional diffusion models. The conditioning input typically consists of historical time series, and the model learns to reconstruct future values from noise under this context. Recent research [65, 80, 81] in diffusion-based TSF also introduces guidance techniques to control the influence of conditioning signals during generation, enabling more flexible and robust forecasting. The following subsections provide a detailed roadmap of these concepts, covering the mathematical formulation of DDPM and DDIM, and the adaptation of conditional diffusion models with an emphasis on condition guidance mechanisms for TSF.
@@ 2.1 Denoising Diffusion Probabilistic Model (DDPM)
Denoising Diffusion Probabilistic Models (DDPMs) [54] define a generative process by progressively adding and removing noise through a sequence of forward and reverse steps, conditioned on auxiliary information. We denote the original data sample and the conditioning input as $\mathbf{Y}_{0}$ and $\mathbf{c}$, respectively. DDPM constructs a forward diffusion process that corrupts $\mathbf{Y}_{0}$ into a noise distribution and a reverse process that reconstructs $\mathbf{Y}_{0}$ from noise conditioned on $\mathbf{c}$. The following illustrates the details of the forward diffusion process and the reverse denoising process.
Forward Diffusion Process The forward process corrupts the data $\mathbf{Y}_{0}$ using a predefined noise variance schedule $\{\beta_{t}\}_{t=1}^{T}$, where $\beta_{t}$ is the noise variance at diffusion step $t$, and $T$ is the total diffusion steps. The entire forward process forms a Markov chain that is formulated by
@@
where $q(\cdot\mid\cdot)$ denotes the forward process distribution and $\mathbf{Y}_{t}$ is the latent variable at diffusion step $t$. In addition, $\mathbf{Y}_{1:T}$ is a noisy data sequence from diffusion step $1$ to $T$. Meanwhile, $q(\mathbf{Y}_{t}\mid\mathbf{Y}_{t-1})=\mathcal{N}(\cdot;\boldsymbol{\mu},\boldsymbol{\Sigma})$ is a multivariate Gaussian distribution with mean $\boldsymbol{\mu}=\sqrt{1-\beta_{t}}\,\mathbf{Y}_{t-1}$ and covariance $\boldsymbol{\Sigma}=\beta_{t}\mathbf{I}$ (where $\mathbf{I}$ denotes the identity matrix). The latent variable $\mathbf{Y}_{t}$ are sampled from a distribution directly conditioned on $\mathbf{Y}_{0}$:
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$. The distribution of the endpoint $\mathbf{Y}_{T}$ of the forward process obeys a Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ with mean $\mathbf{0}$ and covariance $\mathbf{I}$ (which is represented as $\mathbf{Y}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$).
Reverse Denoising Process Contrary to the forward process, the reverse process reconstructs $\mathbf{Y}_{t-1}$ from $\mathbf{Y}_{t}$ via a transition distribution $p_{\theta}(\cdot\mid\cdot)$ parametrized by $\theta$. This process is formulated as
$p_{\theta}(\mathbf{Y}_{t-1}\mid\mathbf{Y}_{t},\mathbf{c})=\mathcal{N}(\mathbf{Y}_{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{Y}_{t},\mathbf{c},t),\Sigma_{\theta}(\mathbf{Y}_{t},\mathbf{c},t))$ (3)
@@
where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the Gaussian noise added to the clean data sample in the forward process, and $\boldsymbol{\epsilon}_{\theta}(\cdot)$ is the noise prediction network with conditional $\mathbf{c}$ as input. In addition, $\mathbb{E}[\cdot]$ denotes the expectation operation and $\left\|\cdot\right\|^{2}$ represents Euclidean norm function. During inference, the model iteratively applies the conditional reverse transitions starting from $\mathbf{Y}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ as
$\mathbf{Y}_{t-1}\sim p_{\theta}(\mathbf{Y}_{t-1}\mid\mathbf{Y}_{t},\mathbf{c}),\quad t=T,T-1,\dots,1$ (5)
@@ 2.2 Denoising Diffusion Implicit Model (DDIM)
Compared to DDPMs, DDIMs eliminate stochasticity in the reverse process by removing the random noise term at each step, while preserving the marginal distributions of the original diffusion trajectory. The step-skip sampling in DDIMs eliminates the need to iterate over all $T$ time steps in the reverse denoising process. Instead, DDIM samples only on a subset of diffusion steps $t_{s_{1}},t_{s_{2}},\dots,t_{s_{S}}$, where $t_{s_{1}}=T$ and $t_{s_{S}}=0$. Specifically, for sampling step $k$ from $1$ to S1S-1, the DDIMs computes $\mathbf{Y}_{t_{s_{k-1}}}$ from $\mathbf{Y}_{t_{s_{k}}}$ using Eq. (6) and replaces $t$ and t1t-1 with $t_{s_{k}}$ and $t_{s_{k-1}}$, respectively.
Beyond DDIMs, several advanced sampling techniques are proposed to accelerate diffusion models. For instance, Analytic-DPM [82] leverages closed-form solutions to the reverse diffusion ordinary differential equation, enabling direct computation of sample trajectories without iterative simulation and thus improving the speed. DPM-Solver [83] applies high-order numerical solvers to the reverse process, replacing the standard Euler approach with more accurate and efficient integration schemes. These approaches further reduce inference time and improve sample quality by optimizing the computation of the reverse trajectory.
Figure 2: Taxonomy of diffusion-based models for TSF. All approaches are firstly categorized by the condition source, leading to models using historical time series and models using multimodal data. The models in each category are further categorized into different sub-categories. Meanwhile, we group these models by the condition integration process into feature-centric and diffusion-centric strategies, where feature-centric and diffusion-centric approaches are highlighted in the red and yellow background colors, respectively. We also present the section number to illustrate models in different categories.
@@ 2.3 Applying Diffusion Models for TSF
A straightforward approach to applying diffusion models to TSF is to set the original data $\mathbf{Y}_{0}$ as the future time series $\mathbf{x}_{1:H}=(x_{1},x_{2},\ldots,x_{i},\ldots x_{H})$, where $H$ is the prediction horizon and $x_{i}$ denotes the time series value at timestep $i$. Meanwhile, the condition $\mathbf{c}$ in diffusion model equals to the historical time series $\mathbf{x}_{-L+1:0}=(x_{-L+1},x_{-L+2},\ldots,x_{0})$ with $L$ denoting the length of the history time series. In this formulation, the model learns the conditional distribution $q(\mathbf{x}_{1:H}\mid\mathbf{x}_{-L+1:0})$ of future values given historical time series $\mathbf{x}_{-L+1:0}$. The forward diffusion process corrupts the future time series $\mathbf{x}_{1:H}$ with Gaussian noise over $T$ steps, while the reverse process reconstructs $\mathbf{x}_{1:H}$ from noise $\mathbf{x}^{T}_{1:H}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, conditioned on the historical time series $\mathbf{x}_{-L+1:0}$. Specifically, the DDPM for TSF, as illustrated in Fig. 1, defines the forward process as
@@
Meanwhile, the reverse process $p_{\theta}(\mathbf{x}_{1:H}^{t-1}\mid\mathbf{x}_{1:H}^{t},\mathbf{x}_{-L+1:0})$ is parameterized as
$\mathcal{N}\big{(}\mathbf{x}_{1:H}^{t-1};\boldsymbol{\mu}_{\theta}(\mathbf{x}_{1:H}^{t},\mathbf{x}_{-L+1:0},t),\Sigma_{\theta}(\mathbf{x}_{1:H}^{t},\mathbf{x}_{-L+1:0},t)\big{)}$ (10)
@@
The DDIM variant for TSF also modifies the reverse process to a deterministic trajectory, which is formulated as
@@
In the field of conditional generation for vision, guidance techniques are proposed to flexibly control the strength of conditioning and improve sample quality. Recent studies in diffusion-based TSF also adapt guidance mechanisms to incorporate additional conditioning sources without modifying model architectures. Guidance mechanisms provide a way to steer the generation process toward desired properties or constraints. One of these guidance techniques, named classifier guidance, modifies the reverse process by adding the gradient of a classifier [^2] Although TSF denotes this component as a predictor, we use the term “classifier” to align with the original nomenclature. $p_{\text{clf}}(\mathbf{x}_{-L+1:0})$ trained to predict the historical time series $\mathbf{x}_{-L+1:0}$ from the noisy future sequence $\mathbf{x}_{1:H}^{t}$. According to Dhariwal et al. [78], the guidance is simplified to an update to the mean function:
@@
Different from the classifier guidance that requires additional training of a classifier, classifier-free guidance [79] trains the diffusion model to handle both conditional and unconditional generations by randomly dropping the conditioning input during training. In inference, the model combines the conditional and unconditional predictions to update the noise $\boldsymbol{\epsilon}_{\text{guided}}$ by:
@@
where $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{1:H}^{t},\varnothing,t)$ is the unconditional noise prediction, $\varnothing$ represents the historical time series is set to a null tensor, and $s\geq 0$ is the guidance scale controlling the strength of conditioning during sampling ($s=0$ yields the conditional prediction, larger $s$ increases guidance). These mechanisms enable flexible integration of historical and auxiliary information in diffusion-based TSF models.